# EgoHand: Ego-centric Hand Pose Estimation and Gesture Recognition with Head-mounted Millimeter-wave Radar and IMUs

### v1

@@ Abstract
Recent advanced Virtual Reality (VR) headsets, such as the Apple Vision Pro, employ bottom-facing cameras to detect hand gestures and inputs, which offers users significant convenience in VR interactions. However, these bottom-facing cameras can sometimes be inconvenient and pose a risk of unintentionally exposing sensitive information, such as private body parts or personal surroundings. To mitigate these issues, we introduce EgoHand. This system provides an alternative solution by integrating millimeter-wave radar and IMUs for hand gesture recognition, thereby offering users an additional option for gesture interaction that enhances privacy protection. To accurately recognize hand gestures, we devise a two-stage skeleton-based gesture recognition scheme. In the first stage, a novel end-to-end Transformer architecture is employed to estimate the coordinates of hand joints. Subsequently, these estimated joint coordinates are utilized for gesture recognition. Extensive experiments involving 10 subjects show that EgoHand  can detect hand gestures with 90.8% accuracy. Furthermore, EgoHand  demonstrates robust performance across a variety of cross-domain tests, including different users, dominant hands, body postures, and scenes.
@@ 1 Introduction
Figure 1: We introduce EgoHand, a VR interaction gesture recognition system utilizing head-mounted millimeter-wave radar and IMU as an alternative or complement to downward-facing cameras for enhancing personal privacy.
Virtual Reality (VR) technology has presented novel human computer interaction approaches in entertainment [28], education [3], working [5], etc. One of the most critical techniques in VR is gesture recognition. Current gesture recognition solutions mainly include joysticks [27, 1] and cameras [47, 33, 34, 45, 36]. Joysticks can provide accurate hand locomotion, while holding joysticks is an added inconvenience for the user. Cameras provide a device-free solution and have been utilized in latest VR headsets like Apple Vision Pro. However, bottom-facing cameras may expose personal privacy, such as user’s sensitive body parts, family members’ faces, and private residence environments.
To mitigate these issues, we introduce EgoHand. This system provides an alternative solution by integrating millimeter-wave radar and IMUs for hand gesture recognition, thereby offering users an additional option for gesture interaction that enhances privacy protection. Current research in mmWave radar-based gesture recognition typically involves fixing the radar at a stationary position and estimating gestures from a frontal horizontal angle [14, 25, 40, 18, 26, 37, 42, 38], Among these works, the most influential is Google’s Soli project [23, 39]. Besides, several studies have explored using mmWave radars as wearable devices to capture human movements in more dynamic settings. For example, Headar [43] integrates mmWave radar into a smartwatch to detect head nods or shakes, useful for confirmation dialogs. EgoMotion [21] and Argus [11] mount the radar on the head to perform body pose estimation and mesh reconstruction, respectively. In contrast, EgoHand  is the first to mount mmWave radar on the head for ego-centric hand pose estimation and gesture recognition.
This discrepancy renders existing mmWave radar work unsuitable for EgoHand, presenting us with two challenges. Firstly, first-person perspective VR interaction gestures encompass both single-handed and two-handed gestures, requiring automatic differentiation and recognition between them. To address this issue, we adopt a two-stage approach. The first stage utilizes end-to-end Transformer-based architectures [6, 41] to automatically estimate hand keypoint, providing end-to-end solutions capable of directly discerning single-handed or two-handed hand keypoint information. The second stage utilizes hand keypoint information as input and employs a classification approach to recognize gestures. The second challenge stems from irregular fluctuations in mmWave signals due to the movement of the mmWave device itself, as the user’s head moves. These head movements cause the radar to shift position and decrease the gesture recognition accuracy. To address this, we integrate IMUs with the mmWave radar to compensate for the positional changes caused by head motion with the feature fusion strategy.
In EgoHand, raw signals from mmWave radar are transformed into range-Doppler heatmaps and range-angle heatmaps. These, along IMUs data, are inputted separately into the mmWave Radar Encoder and IMU Encoder to learn corresponding radar representation and IMU representation, further concatenated and sequentially passed through the Pose Decoder and Context Decoder. The Pose Decoder generates a set of candidate hand pose features, while the Context Decoder further integrates temporal contextual information to produce a set of hand keypoints through MLP layers. In the training phase, we have synchronized hand movement videos and use Google MediaPipe Hand Landmark SDK to process videos into hand keypoints, later as training supervision. We apply set-based Hungarian matching algorithm [20] to optimize the MLP-generated hand keypoints with supervision. During inference, the MLP-generated hand keypoints with high confidence are fed into ResNet50 [15] to obtain the gesture recognition results in an end-to-end manner.
We conducted evaluation on 10 volunteers across 8 types of gesture interactions in 3 different scenes and postures. The mean per joint position error (MPJPE) for hand pose in EgoHand  was 72.7 mm, with a gesture recognition accuracy of 90.8%. Compared to directly utilizing ResNet50 for gesture recognition, the two-stage network approach improved recognition accuracy by approximately 30%. In cross-person experiments, the 2-shot gesture recognition accuracy was 83.9%, while 2-shot cross-scene experiments achieved 76.9% accuracy, and 2-shot cross-hand experiments reached 77.1% accuracy. These results demonstrate the effectiveness of EgoHand  in both in-domain and cross-domain VR gesture recognition. Our contributions in this work are outlined below:
(1) We introduce EgoHand, the first solution to mount mmWave radar and IMUs on the head for VR interaction gesture recognition. This approach offers advantages in user privacy protection compared to existing bottom-facing cameras, inspiring further research into privacy protection for VR hardware and applications.
(2) We propose a Transformer network for end-to-end generation of hand keypoints. Utilizing keypoints as intermediate features in a two-stage gesture recognition process, the network yields significant improvements over single-stage approaches, with accuracy enhancement up to 30%.
(3) We collect a dataset of VR interaction gestures using head-mounted mmWave radar and IMUs, totaling 26GB. The dataset has been fully open-sourced.
@@ 2 Related Work
2.1 Gesture Recognition with Cameras
In the field of computer vision, a series of methods have been developed for hand gesture recognition, most of which rely on RGB/D cameras that provide rich geometric information [31, 48]. Depth cameras perform excellently in handling complex gestures, such as the effective estimation of gesture joints through the combination of the PointNet model  [13], and Zimmermann’s proposal of the 3D Transformer network to further improve recognition accuracy  [50]. Besides, ModDrop [29] fuses video stream, depth stream, and audio stream to classify specific gestures. Zhou et al.  [49] proposed an adaptive cross-modal learning method, designing unique modal fusion strategies for different gestures to improve recognition accuracy. TMMF  [12] focuses on the single-stage recognition of continuous gestures, emphasizing the importance of temporal information in enhancing the continuity of gesture recognition.
Figure 2: EgoHand  takes the head-mounted millimeter-wave radar signals and IMU data to generate hand keypoints in the first stage, and further estimates the gesture category in the second stage. The camera is used solely for label generation during training and is not involved in inference.
2.2 Gesture Recognition with mmWave Radars
Millimeter-wave radar has significant advantages in spatial resolution and penetration ability, providing a reliable and efficient solution for human gesture recognition and pose estimation [26, 24, 30, 32]. It is noteworthy that recent research has demonstrated the progress of millimeter-wave radar in behavior detection, such as mHomeGes  [24] and mTransSee  [26], which use millimeter-wave signals to achieve real-time arm gesture recognition and environment-independent gesture recognition, respectively. mmASL  [32] extracts frequency features from 60GHz millimeter-wave signals and uses a multitask neural network to recognize american sign language. Pantomime et.al. [30] uses millimeter-wave radar to compute sparse 3D point clouds for gesture recognition in their self-collected dataset.
However, most of these mmWave radar based systems are designed to be fixed installations; adapting them to a user’s free movement in space, and overcoming limitations of viewpoint and location remain challenges, which may limit their application in VR interactions.
@@ 3.1 Data Preprocessing
In our setup, mmWave radar transmits and receives frequency modulated continuous wave (FMCW) signals chirp by chirp (64 chirps per frame, 20 frames per seconds), and mixes the received signals with the transmitted signals to obtain an intermediate frequency (IF) signal [2, 17, 9]. Due to the sparse point cloud in mmWave radar data, typically consisting of only a few points, hand reflections are inadequately represented, and minor posture changes can cause significant variations, leading to inconsistent representations. To address this issue, we adopted a preprocessing method similar to that in mHomeGes [24] to obtain a richer representation of the radar data.
$\bullet$ Range FFT. We apply range FFT on every IF chirp, which reveals the frequency difference between the received and transmitted signals, to obtain distance heatmaps of reflection objects.
$\bullet$ Doppler FFT. There exists a phase difference between neighboring chirps caused by the object movements, we use Doppler FFT on chirps in one frame along the direction of the phase change caused by the Doppler effect to estimate the object’s velocity, the outputs called range-Doppler heatmaps.
$\bullet$ Angle FFT. There are phase differences between the antennas caused by the spatial location. We use angle FFT is executed on the range-Doppler heatmaps along the receiving antenna dimension to acquire range-angle heatmaps, which responds to the distance and angle information of the reflection objects.
We concatenate range-Doppler heatmaps and range-angle heatmaps along the vertical axis for the deep network input, as shown in Fig. 3. In contrast, we do not preprocess IMU series.
@@ 3.2 Deep Network Design
We use an 15.85 mW low-power mmWave radar to capture hand movements. However, the radar signal also contains noise from head movements. To compensate for this, we attach an IMU to the mmWave radar to record its motion. This creates a multimodal fusion problem, and recent research shows that Transformer architectures perform well in multimodal learning [7, 22]. Therefore, we use a Transformer-based structure for representation learning.
Another design choice arose from experimental results: directly predicting gesture classes led to poor performance, with accuracy around 65%. We therefore opted to first generate an intermediate representation as a transition step, then use it to classify gestures. Research using cameras [8, 16] has shown that hand skeletons serve as a promising and interpretable intermediate representation, so we adopted this approach.
@@
The third design consideration is handling gestures involving different numbers of hands: 0, 1, or 2 hands may be present during interactions. Unlike OCHID-Fi [46], which assumes a fixed output of keypoints for one hand, we require an approach that can adapt to varying numbers of hands. This is similar to the object detection task, where the model must flexibly output coordinates for a variable number of targets. Therefore, we adapt DETR [6], an end-to-end object detection framework, to regress the coordinates of hand keypoints.
Based on the above considerations, we propose the framework shown in Fig. 2 and the network structure illustrated in Fig. 4. More detailed operations will be introduced next.
@@
Figure 4: mmVR deep network consists of four main novel components. (a) mmWave Radar Encoder, (b) IMU Encoder, (c) Pose Decoder, and (d) Context Decoder.
@@
$\bullet$ mmWave Radar Encoder. This encoder integrates six encoder blocks to process mmWave radar streams for mmWave radar embeddings. Each block comprises a multi-head self-attention module and a Feed-Forward Network (FFN), which are fundamental components of the Transformer architecture  [35].
@@
$\bullet$ IMU Encoder. This part of the architecture is similar to the mmWave Radar Encoder, which is used for IMU data embeddings. The encoded IMU embeddings are concatenated with the mmWave radar embeddings for further Pose Decoder and Context Decoder.
@@
$\bullet$ Pose Decoder. The structure of the decoder layer is the basic module of the DETR [6] decoder, and we stacked three such layers. One Pose Decoder block contains one multi-head self-attention module and one multi-head cross-attention module. The multi-head cross-attention module compute attention matrix between Pose Queries ($Q_{pose}=[q_{1},q_{2},\ldots,q_{C}]\in\mathbb{R}^{C\times D}$) and the fused embeddings ($E_{fused}$) from mmWave Radar Encoder and IMU Encoder. Pose Decoder outputs $X_{pose}\in\mathbb{R}^{C\times D}$, representing features of $C$ hand candidates in $D$ dimension. The process of Pose Decoder can be formulated as follows:
@@
$\bullet$ Context Decoder. One Context Decoder block includes one multi-head self-attention module and two multi-head cross-attention modules, and we stack 30 such blocks in the Context Decoder. Context Decoder learns the globally temporal relationships among different frames to refine the hand keypoint estimation. Specifically, Context Decoder utilizes the multi-head cross-attention to compute cross-attention matrix between the current query and previous pose features from Pose Decoder, iteratively. At last, two Multi-Layer Perceptrons (MLPs) are respectively to generate the keypoints of the hand candidates and the corresponding confidence scores indicating the authenticity of these hand candidates. The process of Context Decoder can be formulated as follows:
@@
where $Q_{context}\in\mathbb{R}^{C\times D}$ represents $C$ learnable context queries; $H_{set}\in\mathbb{R}^{C\times n\times K\times 3}$ represents K 3D-keypoints of C hand candidates of $n$ frames. $H_{score}\in\mathbb{R}^{C\times n}$ represents confidence scores of these hand candidates in $n$ frames. In our settings, $C=100$, $K=21$, and $n=30$.
@@
where $H_{gt}$ denotes the ground-truth hand keypoint coordinates. $H_{set}$ is selected up to two hand candidates based on $H_{score}$. If there are two hands, their keypoints are concatenated and fed into the recognition network. If there is only one hand, it is duplicated before concatenation, ensuring that the recognition network has a fixed input size.
Further, we apply a cross-entropy function to compute the gesture recognition loss.
$L_{cls}=\text{CrossEntropy}(y,\hat{y})$ (4)
where $y$ and $\hat{y}$ represent the gesture label and recognition confidence score, respectively.
The total loss to optimize the deep network is a weighted combination of hand keypoint regression loss and the gesture regression loss.
$Loss=L_{kpt}+\lambda L_{cls}$ (5)
where $\lambda$ is a weight parameter balancing the hand keypoint regression and gesture recognition.
$\bullet$ Training Details. We employed the Adam optimizer [19] to train deep network on a NVIDIA 3090 GPU. The batch size was 32. The training was carried out 500 epochs. The initial learning rate was 0.001 and was halved every 100 epochs. At the start of each epoch, the training samples were randomly shuffled. To balance the keypoint regression and gesture recognition, we experimentally set $\lambda=1$ in Equation. 5.

### v2

# mmEgoHand: Egocentric Hand Pose Estimation and Gesture Recognition with Head-mounted Millimeter-wave Radar and IMU

@@ Abstract
Recent advancements in millimeter-wave (mmWave) radar have demonstrated its potential for human action recognition and pose estimation, offering privacy-preserving advantages over conventional cameras while maintaining occlusion robustness, with promising applications in human-computer interaction and wellness care. However, existing mmWave systems typically employ fixed-position configurations, restricting user mobility to predefined zones and limiting practical deployment scenarios. We introduce mmEgoHand, a head-mounted egocentric system for hand pose estimation to support applications such as gesture recognition, VR interaction, skill digitization and assessment, and robotic teleoperation. mmEgoHand  synergistically integrates mmWave radar with inertial measurement units (IMUs) to enable dynamic perception. The IMUs actively compensate for radar interference induced by head movements, while our novel end-to-end Transformer architecture simultaneously estimates 3D hand keypoint coordinates through multi-modal sensor fusion. This dual-modality framework achieves spatial-temporal alignment of mmWave heatmaps with IMUs, overcoming viewpoint instability inherent in egocentric sensing scenarios. We further demonstrate that intermediate hand pose representations substantially improve performance in downstream task, e.g., VR gesture recognition. Extensive evaluations with 10 subjects performing 8 gestures across 3 distinct postures- standing, sitting, lying - achieve 90.8% recognition accuracy, outperforming state-of-the-art solutions by a large margin. Dataset and code are available at https://github.com/WhisperYi/mmVR.
Index Terms:
Human sensing, Hand pose estimation, Gesture recognition, Millimeter-wave radar, IMUs, Human-computer interaction
@@ 1 Introduction
Figure 1: We present mmEgoHand, an interaction gesture recognition system utilizing head-mounted millimeter-wave radar and IMUs as an alternative or complement to downward-facing cameras for enhancing personal privacy.
Human activity recognition and pose estimation are foundational technologies that support a wide range of applications, including human behavior analysis, wellness care, human-computer interaction, and gaming. Traditionally, these tasks have relied on camera-based solutions [1, 2]. More recently, wireless sensing alternatives such as Wi-Fi [3] and custom-built radar systems [4] have emerged to address the limitations of visual methods, offering improved privacy and robustness to occlusion. However, Wi-Fi-based approaches are sensitive to environmental variations and subject orientation [5, 6], while custom radar systems face challenges in adoption due to hardware accessibility. These constraints have spurred interest in commercially available millimeter-wave (mmWave) radar systems, such as the Texas Instruments (TI) IWR/AWR series, which offer a compelling trade-off between deployment practicality and high-resolution motion sensing for human activity recognition and pose estimation.
The evolution of mmWave radar-based human sensing has progressed from coarse activity recognition [7, 8, 9, 10, 11, 12, 13] to fine-grained pose estimation [14, 15, 16, 17, 18, 19, 20, 21, 22]. Most existing systems employ frontal, static deployments, requiring users to stay within a constrained sensing volume—an approach that limits applicability in dynamic environments. Recent advances in egocentric sensing, such as mmEgo [23] and Argus [24], have begun to overcome this limitation by leveraging head-mounted radar to capture body-reflected signals for pose estimation. These systems enable full-body tracking while preserving human movement, shifting mmWave sensing from environment-anchored perception to mobile, user-centric motion capture.
In this work, we introduce mmEgoHand, an egocentric hand pose estimation system. While prior mmWave methods primarily focus on full-body tracking, our system targets fine-grained hand articulation to support applications such as gesture recognition, VR interaction, skill digitization and assessment, and robotic teleoperation. The design of mmEgoHand  is guided by three key considerations: (1) Existing frontal-view systems [25, 26, 27] typically estimate single-hand poses, whereas many real-world gestures involve both hands. We thus design a unified framework capable of handling both single- and two-hand configurations. (2) Head-mounted radar introduces ego-motion artifacts, as head movements dynamically alter the radar’s perspective, changing signal characteristics for the same hand gesture. Our system incorporates active motion compensation to disentangle hand motion from head-induced signal variations. (3) Traditional motion capture systems, such as VICON [28], are cumbersome for annotating hand poses. We leverage a lightweight annotation pipeline that ensures high-quality ground truth while significantly simplifying deployment, improving adaptability across application scenarios.
To address the first consideration, we adapt an end-to-end Wi-Fi-based multi-person pose estimation [29]. In the original design, the system automatically detects a variable number of individuals and estimates their body keypoints. Similarly, mmEgoHand  is designed to output a variable number of hands with a set-based Hungarian matching algorithm [30] and estimate keypoints for each detected hand in an end-to-end manner. This modification enables mmEgoHand  to support both one-handed and two-handed interaction tasks, making it more versatile and applicable to a wider range of scenarios compared to existing methods such as mm4Arm [27] and mmHand [26, 25], which only support fixed one-handed interactions. Furthermore, we augment the original architecture with a context decoder that takes a sequence of mmWave heatmaps as input, leveraging temporal information to improve the accuracy of hand pose estimation.
For the second consideration, we draw inspiration from the hardware setup of mmEgo [23], attaching an inertial measurement unit (IMU) near the mmWave radar to capture head motion. The IMU data is temporally synchronized with the mmWave input and fed jointly into the network. By fusing these two modalities, mmEgoHand  can compensate for ego-motion artifacts caused by head movement, leading to more stable and accurate hand pose estimation. To address the third consideration, we employ a lightweight annotation strategy during data collection. A standard web camera is placed next to the subject to record hand movements, and hand keypoints are automatically extracted from the video using Google MediaPipe Hand Landmark SDK. We then perform extensive manual checking and filtering, removing keypoints with poor tracking quality, particularly under fast hand motion conditions, and take the remaining as ground truth annotations for training and evaluation. Similar camera-based automatic annotation schemes have become increasingly popular in prior work [3, 4, 24], offering a practical trade-off between annotation quality and deployment simplicity.
To evaluate mmEgoHand, we recruited 10 volunteers who performed hand movements in three different scenes and under three distinct postures: standing, sitting, and lying. mmEgoHand  achieved a mean per-joint position error (MPJPE) of 72.73mm for hand pose estimation. In comparison, using mmWave radar alone resulted in an MPJPE of 96.42mm, highlighting the effectiveness of incorporating IMU data. Furthermore, we used the estimated hand poses from mmEgoHand  as intermediate representations for a downstream hand gesture recognition task. mmEgoHand  achieved an accuracy of 90.80%, substantially outperforming state-of-the-art approaches such as mGesNet [7] (81.34%), mSeeNet [9] (84.60%), and mmGesture [31] (68.66%). These results highlight that mmEgoHand  not only enhances hand pose estimation, but also provides more reliable and informative features for downstream applications. Our contributions are summarized as follows:
(1) We introduce mmEgoHand, the first egocentric system capable of human hand pose estimation and gesture recognition (see Table I for a comparison with prior work).
(2) We propose a novel dual-decoder architecture that simultaneously addresses spatial information and temporal coherence, yielding significant improvements over state-of-the-art approaches
(3) We collect and open-source a large-scale dataset of VR interaction gestures captured using head-mounted mmWave radar and IMU, totaling 26GB, to support future research in human sensing. Dataset and code are available at https://github.com/WhisperYi/mmVR
TABLE I: mmEgoHand  is the first egocentric system capable of simultaneously capturing both one-hand and two-hand poses using mmWave radar.
Solutions Radar Sensing View Body Part
mmMesh [14] AWR1843 Frontal Body
SynMotion [15] IWR1443 Frontal Body
mmGPE [16] AWR1843 Frontal Body
M4esh [17] AWR1843 Frontal Body
mmBody [18] Arbe Phoenix Frontal Body
m3Track [19] IWR1443 Frontal Body
HUPR [20] IWR1843 Frontal Body
Mi-Mesh [21] AWR1443 Frontal Body
SUPER [22] IWR6843 Frontal Upper
mm4Arm [27] IWR6843 Frontal One hand
mmHand [26] AWR1443 Frontal One hand
mmHand [25] IWR1843 Frontal One hand
mmEgo [23] IWR6843 Egocentric Body
Argus [24] BGT60TR13C $\times$ 2 Egocentric Body
mmEgoHand (Ours) IWR6843 Egocentric 1 or 2 Hands
@@ 2 Related Work
2.1 Hand Perception with Cameras
In the field of computer vision, a series of methods have been developed for hand pose estimation and gesture recognition, most of which rely on RGB/D cameras that provide rich geometric information [32, 33]. Depth cameras perform excellently in handling complex gestures, such as the effective estimation of gesture joints through the combination of the PointNet model  [34], and Zimmermann’s proposal of the 3D Transformer network to further improve recognition accuracy  [35]. Besides, ModDrop [36] fuses video stream, depth stream, and audio stream to classify specific gestures. Zhou et al.  [37] proposed an adaptive cross-modal learning method, designing unique modal fusion strategies for different gestures to improve recognition accuracy. TMMF  [38] focuses on the single-stage recognition of continuous gestures, emphasizing the importance of temporal information in enhancing the continuity of gesture recognition.
However, these vision-based methods inherently depend on capturing detailed visual information of users’ hands and surrounding environments, which raises privacy concerns in sensitive applications and shared spaces. In addition, they may suffer from occlusions in egocentric scenarios, where hands frequently self-occlude or interact with objects (e.g., VR controllers).
2.2 Hand Perception with mmWave Radars
In the field of radio frequency (RF) sensing, studies have shown that Wi-Fi and customized radars are capable of estimating the poses of multiple people [3, 4, 29, 39], but the performance in terms of stability and accuracy under complex environments is still lacking, and often requires additional hardware such as increased antenna count and device numbers to enhance performance. In contrast, commercial millimeter-wave radar has significant advantages in spatial resolution and penetration ability, providing a reliable and efficient solution for human gesture recognition and pose estimation [9, 7, 40, 41]. It is noteworthy that recent research has demonstrated the progress of millimeter-wave radar in behavior detection, such as mHomeGes [7] and mTransSee [9], which use millimeter-wave signals to achieve real-time arm gesture recognition and environment-independent gesture recognition, respectively. mmASL [41] extracts frequency features from 60GHz millimeter-wave signals and uses a multitask neural network to recognize American Sign Language. Pantomime et al. [40] uses millimeter-wave radar to compute sparse 3D point clouds for gesture recognition in their self-collected dataset.
It can be observed in Table. I that most existing methods rely on radars fixed in a static position, constraining users to operate strictly within the radar’s frontal field of view. To overcome this limitation, mmEgo [23] and Argus [24] propose mounting radars on the user’s head to enable egocentric body tracking, allowing more natural and unrestricted movements. Building on this design, our work extends egocentric millimeter-wave sensing to fine-grained hand pose estimation and hand gesture recognition, broadening the scope of egocentric radar-based interaction. Moreover, to the best of our knowledge, our system is the first egocentric mmWave-based solution capable of supporting both one-handed and two-handed interactions, significantly expanding the range of applicable interaction scenarios.
@@ 3.1 Data Preprocessing
In our setup, we use a TI IWR6843 mmWave radar, a 1.75 W low-power radar, to capture hand movements. mmWave radar transmits and receives frequency modulated continuous wave (FMCW) signals chirp by chirp (64 chirps per frame, 20 frames per seconds), and mixes the received signals with the transmitted signals to obtain an intermediate frequency (IF) signal [42, 43, 21]. Due to the sparse point cloud in mmWave radar data, typically consisting of only a few points, hand reflections are inadequately represented, and minor posture changes can cause significant variations, leading to inconsistent representations. To address this issue, we adopted a preprocessing method similar to that in mHomeGes [7] to obtain a richer representation of the radar data.
@@
$\bullet$ Range FFT. We apply range FFT on every IF chirp, which reveals the frequency difference between the received and transmitted signals, to obtain distance heatmaps of reflection objects. To reduce interference from reflections in the surrounding environment, we retain only the signals within 2 meters after the range FFT, as the distance from the user’s head-mounted mmWave radar to the hands is typically less than 2 meters for most body types.
$\bullet$ Doppler FFT. There exists a phase difference between neighboring chirps caused by the object movements, we use apply Doppler FFT on chirps in one frame along the direction of the phase change caused by the Doppler effect to estimate the object’s velocity, the outputs called range-Doppler heatmaps. The velocity information is illustrated at the horizontal axis and the distance information at the vertical axis in Fig. 2.
The Range FFT and Doppler FFT can be formulated as follows:
$S_{rd}=\mathcal{F}_{chirp}{(\mathcal{F}_{sampling}{(\text{IF})})}$ (1)
where $S_{rd}$ represents the range-Doppler heatmaps; $\mathcal{F}_{sampling},\mathcal{F}_{chirp}$ signifies the FFT pertaining to the sampling point dimension and chirp dimension, respectively.
Figure 3: mmEgoHand  takes the head-mounted millimeter-wave radar signals and IMU data to generate hand keypoints. The camera is used solely for label generation during training and is not involved in inference.
$\bullet$ Angle FFT. There are phase differences between the antennas caused by the spatial location. We use angle FFT is executed on the range-Doppler heatmaps along the receiving antenna dimension to acquire range-angle heatmaps, which responds to the distance and angle information of the reflection objects. The angle information is illustrated at the horizontal axis and the distance information at the vertical axis in Fig. 2. The process can be formulated as follows:
$S_{ra}=\mathcal{F}_{rx}{(S_{rd})}$ (2)
where $S_{ra}$ denotes the range-angle heatmaps, and $\mathcal{F}_{rx}$ refer to the FFT operation along the receiving antenna dimension.
We concatenate range-Doppler heatmaps and range-angle heatmaps along the vertical axis for the deep network input, as shown in Fig. 2.
@@ 3.2 Deep Network Design
We use TI IWR6843 mmWave radar to capture hand movements. However, the radar signal also contains noise from head movements. To compensate for this, we attach an IMU to the mmWave radar to record its motion. This creates a multimodal fusion problem, and recent research shows that Transformer architectures perform well in multimodal learning [44, 45]. Therefore, we use a Transformer-based structure for representation learning.
@@
Another consideration is handling poses involving different numbers of hands: 0, 1, or 2 hands may be present during interactions. Unlike mmHand [25, 26], which assumes a fixed output of keypoints for one hand, our system must dynamically adapt to variable hand counts. This is similar to the multi-person pose estimation task, where the model must flexibly output coordinates for a variable number of people. To address this, we adapt the Person-in-WiFi 3D framework [29], an end-to-end solution for multi-person pose estimation, to regress hand keypoints for multiple hands. Guided by these insights, we propose the framework shown in Fig. 3 and the network structure illustrated in Fig. 4.
$\bullet$ Input and Output. We take a 2-second sequence of mmWave radar and IMU signals as input and output the corresponding 3D hand pose sequence. The mmWave input is represented as $x_{mm}\in\mathbb{R}^{30\times 256\times 128}$, and the IMU input as $x_{imu}\in\mathbb{R}^{30\times 2\times 3}$, where 30 is the number of sampled frames. The output is a 3D keypoint sequence of shape $\mathbb{R}^{30\times h\times K\times 3}$, where $h$ is the number of hands, and $K$ is the number of keypoints per hand. To align the modalities temporally, we divide the 2-second window into 30 uniform patches. Each mmWave patch is flattened into a vector of shape $\mathbb{R}^{32,768\times 1}$ (i.e., $256\times 128$), and each IMU patch—consisting of one sample across 6 channels (3-axis acceleration and 3-axis angular velocity)—is flattened into a vector of shape $\mathbb{R}^{6\times 1}$. These patch vectors are then independently fed into an mmWave encoder and an IMU encoder to extract modality-specific temporal features for subsequent fusion.
@@
Figure 4: mmEgoHand  deep network consists of four main novel components. (a) mmWave Radar Encoder, (b) IMU Encoder, (c) Pose Decoder, and (d) Context Decoder.
@@
$\bullet$ mmWave Radar Encoder. This encoder integrates six encoder blocks to process mmWave radar streams for mmWave radar embeddings. Each block comprises a multi-head self-attention module and a Feed-Forward Network (FFN), which are fundamental components of the Transformer architecture [46], shown in Fig. 4 (a).
@@
$\bullet$ IMU Encoder. This encoder is similar to the mmWave Radar Encoder, which is used for IMU data embeddings. The encoded IMU embeddings are concatenated with the mmWave radar embeddings for further Pose Decoder and Context Decoder, shown in Fig. 4 (b).
@@
$\bullet$ Pose Decoder. The structure of the decoder layer is the basic module of the Person-in-WiFi 3D [29] decoder, and we stacked three such layers. As shown in Fig. 4 (c), one Pose Decoder block contains one multi-head self-attention module and one multi-head cross-attention module. The multi-head cross-attention module compute attention matrix between Pose Queries ($Q_{pose}=[q_{1},q_{2},\ldots,q_{C}]\in\mathbb{R}^{C\times D}$) and the fused embeddings ($E_{fused}$) from mmWave Radar Encoder and IMU Encoder. Pose Decoder outputs $X_{pose}\in\mathbb{R}^{C\times D}$, representing features of $C$ hand candidates in $D$ dimension. The process of Pose Decoder can be formulated as follows:
@@
where $E_{fused}$ represents the fused features from mmWave radar embeddings and IMU embeddings.
$\bullet$ Context Decoder. As shown in Fig. 4 (d), one Context Decoder block includes one multi-head self-attention module and two multi-head cross-attention modules, and we stack 30 such blocks in the Context Decoder. Context Decoder learns the globally temporal relationships among different frames to refine the hand keypoint estimation. Specifically, Context Decoder utilizes the multi-head cross-attention to compute a cross-attention matrix between the current query and previous pose features from Pose Decoder, iteratively. At last, two Multi-Layer Perceptrons (MLPs) are respectively to generate the keypoints of the hand candidates and the corresponding confidence scores indicating the authenticity of these hand candidates. The process of Context Decoder can be formulated as follows:
@@
where $Q_{context}\in\mathbb{R}^{C\times D}$ represents $C$ learnable context queries; $H_{set}\in\mathbb{R}^{C\times n\times K\times 3}$ represents K 3D-keypoints of C hand candidates of $n$ frames. $H_{score}\in\mathbb{R}^{C\times n}$ represents confidence scores of these hand candidates in $n$ frames. In our settings, $C=100$, $K=21$, and $n=30$. The Context Decoder consists of 30 Transformer blocks, each responsible for predicting the pose at a single time step. Each block takes as input the current frame along with the refined representation from the previous frame, enabling cross-frame temporal modeling through cross-attention.
This dual-decoder design explicitly decomposes the egocentric hand tracking challenge: the Pose Decoder specializes in spatial joint localization by resolving per-frame ambiguities through cross-attention between pose queries $Q_{pose}$ and sensor embeddings $E_{fused}$; conversely, the Context Decoder captures temporal gesture kinematics via iterative feature refinement across frames, enabling robust recognition of dynamic motions. This separation aligns with the observation that spatial precision and temporal coherence require specialized processing under head-motion interference.
@@
where $H_{gt}$ denotes the ground-truth hand keypoint coordinates. $H_{set}$ is selected up to two hand candidates based on $H_{score}$.
$\bullet$ Implementation Details. We employed the Adam optimizer [47] to train a deep network on an NVIDIA 3090 GPU. The batch size was 32. The training was carried out for 500 epochs. The initial learning rate was 0.001 and was halved every 100 epochs. At the start of each epoch, the training samples were randomly shuffled.
The resulting model contains 55.42M parameters, with an inference GPU memory of 657 MB and an end-to-end inference time of 42.03ms for 30 frames on the RTX 3090 GPU. These results demonstrate that our framework already achieves real-time performance.
$\bullet$ Downstream gesture recognition task. The output of mmEgoHand  can be seamlessly applied to downstream tasks such as gesture recognition. Given a sequence of predicted hand poses $H_{set}$, we feed it into a gesture classification network (e.g., a GNN or ResNet) to identify the performed gesture. To ensure a consistent input size for the recognition model, we handle single- and dual-hand scenarios differently: if two hands are present, their keypoints are concatenated directly; if only one hand is detected, its keypoints are duplicated before concatenation. This design allows the recognition network to operate on a fixed-size input regardless of the number of hands.